# EvoLattice: Persistent Internal-Population Evolution through Multi-Alternative Quality-Diversity Graph Representations for LLM-Guided Program Discovery

### v1

@@ 1 Introduction
Even when populations or genealogical trees are maintained, the representation of each candidate remains internally single-path. None of these approaches represents multiple alternatives for the same operation inside the candidate itself. As a result, the system cannot preserve multiple viable variants of the same subroutine, recombine high-performing components from different individuals, or accumulate a persistent library of reusable micro-operators. The absence of internal quality-diversity forces these methods to rediscover effective patterns repeatedly and prevents them from exploiting the combinatorial richness that multi-alternative representations enable.
@@
Existing systems give the LLM a view of the current program or agent and sometimes a textual summary of ancestry, but the model never sees a structured, quantitative representation of how past alternatives performed. It cannot tell which subroutines consistently improved performance, which showed promise in specific contexts, or which combinations exhibited synergistic effects. It has no access to per-operator contributions, no path-level decomposition, and no principled way to assess how new proposals should align with historical evidence. This informational bottleneck forces the LLM to mutate based on limited context, reducing its ability to perform guided exploration, crossover, specialization, or targeted refinement.
Notably, the resulting evolutionary dynamics resemble those of QD methods such as MAP-Elites (fontaine2020mapelites), where progress arises from preserving and refining a diverse repertoire of partial solutions rather than optimizing population averages, but in existing LLM-driven systems this behavior emerges only implicitly and transiently, without structural support or persistent internal diversity.
@@
We introduce EvoLattice, a representational shift that enables LLM-driven evolution to operate over a structured population encoded inside a single artifact. Instead of storing only one implementation per component, an EvoLattice node holds multiple alternatives, each representing a different micro-operator, subroutine, or agent-level behavior. All alternatives remain active unless structurally pruned, and all possible combinations of alternatives across the graph are executable paths. This creates an exponentially rich landscape of candidate programs or agents, all encoded compactly inside one unified representation.
@@
A defining feature of EvoLattice is its alternative-level performance statistics. Every alternative at every node is evaluated across all valid paths that include it, with caching ensuring computational efficiency. This produces per-alternative estimates of contribution to global performance—including mean effect, variability, and best-case influence. These statistics are presented directly to the LLM alongside the structural graph, providing a unified, quantitative view of the entire evolutionary landscape. The LLM can therefore reason about quality–diversity explicitly, performing targeted mutation, crossover, specialization, or pruning based on empirical signal rather than heuristic guesswork.
After each mutation, a deterministic self-repair pipeline restores acyclicity, removes dangling dependencies, prunes empty nodes, and ensures reachability. This shifts responsibility for structural correctness away from the LLM, allowing the model to focus on high-level reasoning rather than syntactic fragility. EvoLattice thus transforms the evolutionary loop into a multi-path, memory-preserving, statistically informed, and self-repairing process—unifying program and agent evolution within the same structural framework.
@@ 3 Methodology
The mutation is driven by an LLM acting on a structured representation of the entire graph. Before each mutation step, the system automatically generates a set of structured hypotheses derived from the current snapshot with its alternative-level statistics, and from the full task implementation. These hypotheses provide domain-aware reasoning that is embedded directly into the mutation prompt, giving the LLM a richer context for targeted, high-quality modifications. In addition to statistics, the mutation prompt also includes a lightweight structural-diff signal: a unified diff between the previous and mutated EvoLattice, summarizing only the components modified by the last mutation. This diff acts as a compact behavioral trace that helps the LLM understand how its prior hypotheses actually changed the structure and whether those changes improved performance, enabling more informed hypothesis refinement. The LLM then receives the set of nodes and alternatives, their statistics, the dependency structure, and the relevant constraints, and may propose deletion of weak alternatives, creation of new alternatives within existing nodes, creation of new nodes (with at least one alternative), or adjustments to dependencies. Unlike diff-based approaches, the LLM does not rewrite an entire program or agent; instead, it performs local, compositional edits that expand or refine the space of alternatives without erasing accumulated memory. Mutations are orchestrated by a two-level prompting architecture: a global system prompt encoding structural rules and a per-iteration mutation prompt containing the current graph, statistics, and generated hypotheses. This separation stabilizes LLM behavior and ensures task-general constraints are never overwritten. To prevent structural corruption, the EvoLattice is equipped with a deterministic self-repair pipeline applied after every mutation. This pipeline enforces acyclicity, removes invalid or unreachable structures, prunes nodes with no alternatives, and ensures no alternative depends on the output node, allowing the LLM to focus exclusively on creative structural proposals.
@@
The full evolutionary step proceeds as: (i) evaluate all paths (or a subset) with memoized execution; (ii) compute alternative-level statistics; (iii) present the entire EvoLattice and statistics to the LLM; (iv) mutate according to LLM proposals; (v) repair the graph to enforce structural invariants; and (vi) accept or reject the mutation based on performance improvement. Because the representation encodes a population internally, each step is both exploratory (new alternatives expand the search space) and exploitative (statistics highlight strong components), naturally supporting quality-diversity without external archives or novelty pressure. EvoLattice extends seamlessly to agent ecosystems. Nodes represent agent roles, tool-use policies, or prompt fragments; alternatives represent different formulations of these behaviors. Execution of a path corresponds to running a complete multi-agent system. Caching accelerates repeated tool interactions, and alternative-level statistics reflect how each sub-policy contributes across many cooperative configurations, and also how specific prompt instructions, tool configurations, or sub-agent behaviors influence global agent performance across all pathways in which they appear.
@@
Diff-based methods maintain exactly one candidate per iteration, and population-based methods maintain $m$ candidates, typically with $m\ll\prod_{i=1}^{n}k_{i}$ where each EvoLattice node $v_{i}$ has $k_{i}$ alternatives. Thus, EvoLattice maintains a representation whose expressive capacity scales multiplicatively, not linearly, with the number of modifications introduced over time. In traditional LLM-driven evolution, a mutation is destructive: the new candidate is unrelated to the previous except through textual ancestry. EvoLattice instead performs monotonic expansion of the search space by increasing the number of alternatives at relevant nodes while preserving existing ones. Deletions may remove weak alternatives, but these are targeted, not structural, and occur only when supported by performance statistics. As a result, the space of candidates never shrinks unless explicitly guided by performance. Self-repair ensures the EvoLattice satisfies acyclicity, consistency, reachability, and non-emptiness constraints. Let $R$ be the repair operator. Then for any LLM mutation $M$, the updated graph $\widetilde{G}_{t+1}=R(M(G_{t}))$ is guaranteed to lie in the valid subspace of graphs. Diff-based methods rely on the LLM to remain within this subspace, which fails unpredictably.
@@ 4 Experiments
These methods use LLMs to generate or evolve a single proxy function: FunSearch-style regeneration, AlphaEvolve-style diff evolution, and ShinkaEvolve-style ensemble sampling. Although mutation strategies differ, all these methods evolve a single-path proxy and overwrite prior candidates.
All experiments are conducted on NAS-Bench-Suite-Zero (Medium), a synthetic Transformer language-model benchmark designed for large-scale zero-shot architecture ranking. Each sampled architecture is instantiated as a Transformer LM with randomized depth, width, MLP ratio, attention configuration, activation function, and positional encoding. Architectures are evaluated without training using sampled cross-entropy loss, spectral activation statistics, and covariance-based statistics. A teacher score is computed as a weighted combination of negative loss, spectral proxy, and covariance proxy. Proxy quality is measured by Spearman rank correlation ($\rho$) between proxy scores and teacher scores across architectures. All methods are evaluated using a two-phase protocol:

### v2

@@ 1 Introduction
Even when populations or genealogical trees are maintained, the representation of each candidate remains internally single-path. None of these approaches represents multiple alternatives for the same operation inside the candidate itself. As a result, the system cannot preserve multiple viable variants of the same subroutine, recombine high-performing components from different individuals, or accumulate a persistent library of reusable micro-operators. The absence of internal quality-diversity forces these methods to rediscover effective patterns repeatedly and prevents them from exploiting the combinatorial richness that multi-alternative representations enable. Existing systems give the LLM a view of the current program or agent and sometimes a textual summary of ancestry, but the model never sees a structured, quantitative representation of how past alternatives performed. It cannot tell which subroutines consistently improved performance, which showed promise in specific contexts, or which combinations exhibited synergistic effects. It has no access to per-operator contributions, no path-level decomposition, and no principled way to assess how new proposals should align with historical evidence. This informational bottleneck forces the LLM to mutate based on limited context, reducing its guided exploration, crossover, specialization, or targeted refinement ability.
@@
In this work, we introduce EvoLattice, a representational shift that enables LLM-driven evolution to operate over a structured population encoded inside a single artifact. Instead of storing only one implementation per component, an EvoLattice node holds multiple alternatives, each representing a different micro-operator, subroutine, or agent-level behavior. All alternatives remain active unless structurally pruned, and all possible combinations of alternatives across the graph are executable paths. This creates an exponentially rich landscape of candidate programs or agents, all encoded compactly inside one unified representation. Notably, the resulting evolutionary dynamics resemble those of QD methods such as MAP-Elites (fontaine2020mapelites), where progress arises from preserving and refining a diverse repertoire of partial solutions rather than optimizing population averages, but in existing LLM-driven systems this behavior emerges only implicitly and transiently, without structural support or persistent internal diversity.
@@
A defining feature of EvoLattice is its alternative-level performance statistics. Every alternative at every node is evaluated across all valid paths that include it, with caching ensuring computational efficiency. This produces per-alternative estimates of contribution to global performance—including mean effect, variability, and best-case influence. These statistics are presented directly to the LLM alongside the structural graph, providing a unified, quantitative view of the entire evolutionary landscape. The LLM can therefore reason about quality–diversity explicitly, performing targeted mutation, crossover, specialization, or pruning based on empirical signal rather than heuristic guesswork. After each mutation, a deterministic self-repair pipeline restores acyclicity, removes dangling dependencies, prunes empty nodes, and ensures reachability. This shifts responsibility for structural correctness away from the LLM, allowing the model to focus on high-level reasoning rather than syntactic fragility. EvoLattice thus transforms the evolutionary loop into a multi-path, memory-preserving, statistically informed, and self-repairing process—unifying program and agent evolution within the same structural framework.
@@ 3 Methodology
The mutation is driven by an LLM acting on a structured representation of the entire graph. In our implementation, all LLM interactions use openai2025gptoss120bgptoss20bmodel, with temperature set to 0.5 during hypothesis generation to encourage exploratory, domain-aware reasoning, and temperature fixed at 0.0 during mutation steps to ensure deterministic and reproducible structural edits; all generations use a maximum output budget of 65,536 tokens. Before each mutation step, the system automatically generates a set of structured hypotheses derived from the current snapshot with its alternative-level statistics, and from the full task implementation. These hypotheses are embedded directly into the mutation prompt, giving the LLM a richer context for targeted, high-quality modifications. In addition to statistics, the mutation prompt includes a lightweight structural-diff signal: a unified diff between the previous and mutated EvoLattice, summarizing only the components modified by the last mutation. This diff acts as a compact behavioral trace that allows the LLM to assess how its prior hypotheses altered the structure and whether those changes improved performance, enabling informed hypothesis refinement. The LLM then receives the set of nodes and alternatives, their statistics, the dependency structure, and the relevant constraints, and may propose deletion of weak alternatives, creation of new alternatives within existing nodes, creation of new nodes (with at least one alternative), or adjustments to dependencies. Unlike diff-based approaches, the LLM does not rewrite an entire program or agent; instead, it performs local, compositional edits that expand or refine the space of alternatives without erasing accumulated memory. Mutations are orchestrated by a two-level prompting architecture: a global system prompt encoding structural rules and a per-iteration mutation prompt containing the current graph, statistics, and generated hypotheses. This separation stabilizes LLM behavior and ensures task-general constraints are never overwritten. To prevent structural corruption, EvoLattice applies a deterministic self-repair pipeline after every mutation, enforcing acyclicity, removing invalid or unreachable structures, pruning nodes with no alternatives, and ensuring no alternative depends on the output node, thereby allowing the LLM to focus on creative structural proposals.
@@
The full evolutionary step proceeds as: (i) evaluate all paths (or a subset) with memoized execution; (ii) compute alternative-level statistics; (iii) present the entire EvoLattice and statistics to the LLM; (iv) mutate according to LLM proposals; (v) repair the graph to enforce structural invariants; and (vi) accept or reject the mutation based on performance improvement. Because the representation encodes a population internally, each step is both exploratory (new alternatives expand the search space) and exploitative (statistics highlight strong components), naturally supporting quality-diversity without external archives or novelty pressure. EvoLattice extends seamlessly to agent ecosystems. Nodes represent agent roles, tool-use policies, or prompt fragments; alternatives represent different formulations of these behaviors. Execution of a path corresponds to running a complete multi-agent system. Caching accelerates repeated tool interactions, and alternative-level statistics reflect how each sub-policy contributes across many cooperative configurations, and also how specific prompt instructions, tool configurations, or sub-agent behaviors influence global agent performance across pathways they appear.
@@
Diff-based methods maintain exactly one candidate per iteration, and population-based methods maintain $m$ candidates, typically with $m\ll\prod_{i=1}^{n}k_{i}$ where each EvoLattice node $v_{i}$ has $k_{i}$ alternatives. Thus, EvoLattice maintains a representation whose expressive capacity scales multiplicatively, not linearly, with the number of modifications introduced over time. In traditional LLM-driven evolution, a mutation is destructive: the new candidate is unrelated to the previous except through textual ancestry. EvoLattice instead performs monotonic expansion of the search space by increasing the number of alternatives at relevant nodes while preserving existing ones. Deletions may remove weak alternatives, but these are targeted, not structural, and occur only when supported by performance statistics. As a result, the space of candidates never shrinks unless explicitly guided by performance. Self-repair ensures the EvoLattice satisfies acyclicity, consistency, reachability, and non-emptiness constraints. Let $R$ be the repair operator. Then for any LLM mutation $M$, the updated graph $\widetilde{G}_{t+1}=R(M(G_{t}))$ is guaranteed to lie in the valid subspace of graphs, in which diff-based methods rely on the LLM to remain, but fails.
@@ 4 Experiments
These methods use LLMs to generate or evolve a single proxy function: FunSearch-style regeneration, AlphaEvolve-style diff evolution, and ShinkaEvolve-style parent sampling. Although mutation strategies differ, all these methods evolve a single-path proxy and overwrite prior candidates.
All experiments are conducted on NAS-Bench-Suite-Zero (Medium), a synthetic Transformer language-model benchmark designed for large-scale zero-shot architecture ranking. Each sampled architecture is instantiated as a Transformer LM with randomized depth, width, MLP ratio, attention configuration, activation function, and positional encoding. Architectures are evaluated without training using sampled cross-entropy loss, spectral activation statistics, and covariance-based statistics. A teacher score is computed as a weighted combination of negative loss, spectral proxy, and covariance proxy. Proxy quality is measured by Spearman rank correlation ($\rho$) between proxy scores and teacher scores across architectures, evaluated using a two-phase protocol: